# BEASST: Behavioral Entropic Gradient based Adaptive Source Seeking for Mobile Robots

### v1

@@ 8 Room
TABLE I: The complete summarized statistics for all the different scenarios Fig. 6. We show the average total path length and time taken over ten runs for locating sources in different scenarios with different source signal models. The proposed policy (BEASST) is compared against various methods like Infosploration, Behaviorsploration ($\alpha=\{0.8,2.0\}$), and Random Frontiers.
@@
Here, $L_{0}$ is the reference power, $n$ is the decay exponent, $f(.,.)$ models shadowing, and $\epsilon$ models multipath fading. This model closely captures the real-work radio-frequency(RF) signal propagation, accounting for common environmental interference such as obstacles and reflective surfaces, enhancing the realism of our simulation. Alternatively, an exponential decay model [25] can be used directly as a signal strength map (visualized in Fig. 3(f)):
@@
where $\tilde{d}(\bm{\mathbf{x}},\bm{\mathbf{x}}_{s})=-\log\left(\sum_{i=1}^{N}e^{-d_{sp}(\bm{\mathbf{x}},\bm{\mathbf{x}}_{s_{i}})}\right)$ is a soft minimum approximation of the shortest path distance $d_{sp}(\bm{\mathbf{x}},\bm{\mathbf{x}}_{s_{i}})$ between $\bm{\mathbf{x}}\in\mathcal{E}$ and the $i$-th source $\bm{\mathbf{x}}_{s_{i}}$. This model provides an abstraction beneficial for evaluating navigation strategies as it naturally embeds obstacle avoidance. We normalize these signal strength maps as shown above to use them as a surrogate probability density function p(x).
@@
Results: We quantitatively evaluate BEASST’s performance across different environments. Our framework successfully completes source-seeking missions by efficiently balancing source-seeking and exploration, outperforming state-of-the-art methods. Before delving into the results, we want to explain Behaviorsploration, a variation of Infosploration [12] that is used as one of our comparison baselines. In this framework, we replace Shannon entropy with BE as the information gain metric. Our objective is to examine the resulting behavior of robots while SS without any parametric tuning for $\alpha$, thereby illustrating the importance of principled tuning. Essentially, the framework uses frontier-based method and uses BE to find the most perceived informative region using a fixed $\alpha$ to locate the source. We use $\alpha=0.8$ and $2.0$ as these values have been shown in  [13], to produce effective and diverse exploration behaviors. Our set of baselines therefore, includes Infosploration [12], Behaviorsploration (fixed $\alpha=\{0.8,2.0\}$), and Random Frontier Based Exploration (purely random frontier selection when no signal is detected). Each environment was tested over ten trials per scenario, varying source placements, robot initial positions, and number of sources. We measured total time taken and total path length. Table I shows that in environments with consistently detectable signals (e.g., 8-room), BEASST outperforms Infosploration by reaching sources 30 to 50 seconds faster across ten runs for different models. This improvement stems from BEASST’s adaptive tuning, enabling aggressive steps with reliable source signals. We set $\alpha_{max}=2.5$ and $\alpha_{min}=0.5$ in these experiments to prevent gradient overshoot.
Furthermore, in environments with intermittently available source signals (32-room, DARPA SubT), BEASST significantly outperformed all other methods by effectively balancing exploration and source seeking. Figure 7 and Table I quantitatively highlight BEASST’s advantages. In the SubT environment, our framework achieved an approximate 15% reduction in total path length and located the source 180 to 200 seconds faster. This improvement is primarily due to the dynamic and adaptive reweighing of uncertainty using our adaptive tuning mechanism: quick, decisive movements in high-confidence regions and cautious, exploratory behavior in low-confidence regions. Similarly, in the 32-room environment, BEASST also demonstrated a significant reduction in time taken to locate the source and total path length.
@@ VII Conclusion and Future Work
This work presented a novel gradient-based algorithm for robotic source seeking that dynamically adapted trajectory based on signal reliability using Behavioral Entropy with Prelec’s probability weighting function. The approach established theoretical convergence guarantees under unimodal signals, provided practical stability under bounded disturbances, and introduced adaptive parameter tuning for autonomous exploration-exploitation transitions integrated with Gaussian Process-based exploration. Experimental validation across DARPA SubT and multi-room scenarios demonstrated that BEASST consistently outperformed state-of-the-art methods, achieving 180-200 seconds faster convergence with reduced path length through intelligent uncertainty-driven navigation. Future work will directly address measurement noise effects through robust estimation and control design, incorporate robot pose uncertainty into unified localization-mapping-seeking objectives, investigate performance under non-unimodal signal distributions, and validate the approach with physical robots in outdoor environments.

### v2

@@ 8 Room
TABLE I: The complete summarized statistics for all the different scenarios Fig. 6. We show the average total path length and time taken over ten runs for locating sources in different scenarios with different source signal models. The proposed policy (BEASST) and its noisy variant are compared against various methods like Infosploration, Behaviorsploration ($\alpha=\{0.8,2.0\}$), and Random Frontiers.
@@
Here, $L_{0}$ is the reference power, $n$ is the decay exponent, $f(.,.)$ models shadowing, and $\epsilon$ models multipath fading. This model closely captures the real-world radio-frequency(RF) signal propagation, accounting for common environmental interference such as obstacles and reflective surfaces, enhancing the realism of our simulation. Alternatively, an exponential decay model [26] can be used directly as a signal strength map (visualized in Fig. 3(f)):
@@
where $\tilde{d}(\bm{\mathbf{x}},\bm{\mathbf{x}}_{s})=-\log\left(\sum_{i=1}^{N}e^{-d_{sp}(\bm{\mathbf{x}},\bm{\mathbf{x}}_{s_{i}})}\right)$ is a soft minimum approximation of the shortest path distance $d_{sp}(\bm{\mathbf{x}},\bm{\mathbf{x}}_{s_{i}})$ between $\bm{\mathbf{x}}\in\mathcal{E}$ and the $i$-th source $\bm{\mathbf{x}}_{s_{i}}$. This model provides an abstraction beneficial for evaluating navigation strategies as it naturally embeds obstacle avoidance. We normalize these signal strength maps as shown above to use them as a surrogate probability density function p(x). The BEASST-generated behavioral gradient field provide high-level waypoints, which are tracked using an MPPI [27] controller with the system model of our Clearpath Warthog platform, enforcing non-holonomic kinematics, acceleration limits, and obstacle avoidance costs.
@@
Results: We quantitatively evaluate BEASST’s performance across different environments. Our framework successfully completes source-seeking missions by efficiently balancing source-seeking and exploration, outperforming state-of-the-art methods. Before delving into the results, we want to explain Behaviorsploration, a variation of Infosploration [13] that is used as one of our comparison baselines. In this framework, we replace Shannon entropy with BE as the information gain metric. Our objective is to examine the resulting behavior of robots while SS without any parametric tuning for $\alpha$, thereby illustrating the importance of principled tuning. Essentially, the framework uses frontier-based method and uses BE to find the most perceived informative region using a fixed $\alpha$ to locate the source. We use $\alpha=0.8$ and $2.0$ as these values have been shown in  [14], to produce effective and diverse exploration behaviors. Our set of baselines therefore, includes Infosploration [13], Behaviorsploration (fixed $\alpha=\{0.8,2.0\}$), and Random Frontier Based Exploration (purely random frontier selection when no signal is detected). Each environment was tested over ten trials per scenario, varying source placements, robot initial positions, and number of sources. We measured total time taken and total path length. Table I shows that in environments with consistently detectable signals (e.g., 8-room), BEASST outperforms Infosploration by reaching sources 30 to 50 seconds faster across ten runs for different models. This improvement stems from BEASST’s adaptive tuning, enabling aggressive steps with reliable source signals. We set $\alpha_{max}=2.5$ and $\alpha_{min}=0.5$ in these experiments to prevent gradient overshoot. Furthermore, in environments with intermittently available source signals (32-room, DARPA SubT), BEASST significantly outperformed all other methods by effectively balancing exploration and source seeking. Figure 7 and Table I quantitatively highlight BEASST’s advantages. In the SubT environment, our framework achieved an approximate 15% reduction in total path length and located the source 180 to 200 seconds faster. This improvement is primarily due to the dynamic and adaptive reweighing of uncertainty using our adaptive tuning mechanism: quick, decisive movements in high-confidence regions and cautious, exploratory behavior in low-confidence regions. Similarly, in the 32-room environment, BEASST also demonstrated a significant reduction in time taken to locate the source and total path length. We also assessed a noisy gradient implementation in the DARPA SubT Challenge and 32-room environments for two signal models. Specifically, we added zero-mean Gaussian noise, $\bm{\mathbf{\omega}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ where $\|\bm{\mathbf{\omega}}(t)\|\leq 1$, truncated to enforce bounded magnitude in accordance with Lemma 1. We see that compared to the noise-free BEASST, the noisy-gradient variant increases path length and time by $12-15\%$ but in average still outperforms the noise-free benchmarks ($5-11\%$) as shown in the Table I. These results show the effectiveness of our approach and are consistent with Lemma 1: additive gradient errors causes a performance loss but do not break convergence.
@@ VII Conclusion and Future Work
This work presented a novel gradient-based algorithm for robotic source seeking that dynamically adapts trajectory based on signal reliability using Behavioral Entropy with Prelec’s probability weighting function. The approach established theoretical convergence guarantees under unimodal signals, provided practical stability under bounded disturbances, and introduced adaptive parameter tuning for autonomous exploration-exploitation transitions integrated with Gaussian Process-based exploration. Experimental validation across DARPA SubT and multi-room scenarios demonstrated that BEASST consistently outperformed state-of-the-art methods, achieving 180-200 seconds faster convergence with reduced path length through intelligent uncertainty-driven navigation. Future work will address hardware experiments incorporating source measurement noise through robust estimation and control design, incorporate robot pose uncertainty into unified localization-mapping-seeking objectives, and extend the framework to multi-agent scenarios.
Acknowledgments
This work was developed by Donipolo Ghimire during a 2024 summer internship at the U.S. Army Combat Capabilities Development Command Army Research Laboratory (ARL), Adelphi, MD.